OpenStudy (ledah):

Solve for n: -2n + (-7n) + 5 = 59

Parth (parthkohli):

\( \color{Black}{\Rightarrow -2n - 7n = 54}\) \( \color{Black}{\Rightarrow -9n = 54}\) \( \color{Black}{\Rightarrow n = 54 \div -9 }\)

OpenStudy (unklerhaukus):

\[-2n + (-7n) + 5 = 59\]\[−(7+2)n=59−5\]\[n=\frac{59-5}{-(7+2)}=\]

OpenStudy (ledah):

okay I got it. Thanks guys